Price elasticity of demand

find the price elasticity of demand if the demand is given as d= R/2P, where R is the consumers income and P the price
Answer #1

РТ SON 5 7 0 → 90 98 2 = 100- 2p At p=1, as 98 | At p=5. x= 90 When pal consume ₂ x 98 (50-10 surplus 2401 Whin p=5, consumee

answered by: ANURANJAN SARSAM